Abstract

Contents

1 domain Syntax 2 2 domain ConstructorCall 7 3 domain ElaboratedExpression 8 4 License 10

1 domain Syntax

⟨domain SYNTAX Syntax⟩≡

)abbrev domain SYNTAX Syntax
++ Author: Gabriel Dos Reis
++ Date Created: November 10, 2007
++ Date Last Updated: December 05, 2007
++ Description:  This domain provides a simple, general, and arguably
++ complete representation of Spad programs as objects of a term algebra
++ built from ground terms of type boolean, integers, foats, symbols,
++ and strings.  This domain differs from InputForm in that it represents
++ any entity from a Spad program, not just expressions.
++ Related Constructors: Boolean, Integer, Float, symbol, String, SExpression.
++ See Also: SExpression, SetCategory
++ The equality supported by this domain is structural.
++ Fixme: Provide direct support for boolean values, arbritrary
++        precision float point values.
Syntax(): Public == Private where
  Public ==> Join(UnionType, SetCategory) with
    convert: % -> SExpression
      ++ convert(s) returns the s-expression representation of a syntax.
    convert: SExpression -> %
      ++ convert(s) converts an s-expression to syntax.  Note, when `s'
      ++ is not an atom, it is expected that it designates a proper list,
      ++ e.g. a sequence of cons cell ending with nil.
    coerce: Integer -> %
      ++ coerce(i) injects the integer value `i' into the Syntax domain
    coerce: % -> Integer
      ++ coerce(s) extracts and integer value from the syntax `s'
    autoCoerce: % -> Integer
      ++ autoCoerce(s) forcibly extracts an integer value from
      ++ the syntax `s'; no check performed.  To be called only
      ++ at the discretion of the compiler.
    coerce: DoubleFloat -> %
      ++ coerce(f) injects the float value `f' into the Syntax domain
    coerce: % -> DoubleFloat
      ++ coerce(s) extracts a float value from the syntax `s'.
    autoCoerce: % -> DoubleFloat
      ++ autoCoerce(s) forcibly extracts a float value from the syntax `s';
      ++ no check performed.  To be called only at the discretion of
      ++ the compiler

coerce: Symbol -> %
++ coerce(s) injects the symbol `s' into the Syntax domain.
coerce: % -> Symbol
++ coerce(s) extracts a symbol from the syntax `s'.
autoCoerce: % -> Symbol
++ autoCoerce(s) forcibly extracts a symbo from the Syntax
++ domain `s'; no check performed.  To be called only at
++ at the discretion of the compiler.
coerce: String -> %
++ coerce(s) injects the string value `s' into the syntax domain
coerce: % -> String
++ coerce(s) extracts a string value from the syntax `s'.
autoCoerce: % -> String
++ autoCoerce(s) forcibly extracts a string value from
++ the syntax `s'; no check performed.  To be called only at
++ the discretion of the compiler.
buildSyntax: (Symbol, List %) -> %
++ buildSyntax(op, [a1, ..., an]) builds a syntax object for op(a1,...,an).
buildSyntax: (%, List %) -> %
++ buildSyntax(op, [a1, ..., an]) builds a syntax object for op(a1,...,an).
nil?: % -> Boolean
++ nil?(s) is true when `s' is a syntax for the constant nil.
getOperator: % -> Union(Integer, DoubleFloat, Symbol, String, %)
++ getOperator(x) returns the operator, or tag, of the syntax `x'.
++ The return value is itself a syntax if `x' really is an
++ application of a function symbol as opposed to being an
++ atomic ground term.

getOperands: % -> List %
++ getOperands(x) returns the list of operands to the operator in `x'.
compound?: % -> Boolean
++ compound? x is true when not an atomic syntax.
_case: (%, [|Integer|]) -> Boolean
++ x case Integer is true is x really is an Integer

_case: (%, [|DoubleFloat|]) -> Boolean
++ x case DoubleFloat is true is x really is a DoubleFloat
_case: (%, [|Symbol|]) -> Boolean
++ x case Symbol is true is x really is a Symbol
_case: (%, [|String|]) -> Boolean
++ x case String is true is x really is a String
Private ==> SExpression add
rep(x: %): SExpression ==
x pretend SExpression
per(x: SExpression): % ==
x pretend %
x = y ==
EQUAL(x,y)$Lisp @ Boolean
s case Integer ==
integer? rep s
s case DoubleFloat ==
float? rep s
s case String ==
string? rep s
s case Symbol ==
symbol? rep s
convert(x: %): SExpression ==
rep x
convert(x: SExpression): % ==
per x
coerce(i: Integer): % ==
i pretend %
autoCoerce(i: %): Integer ==                -- used for hard coercion
i : Integer
coerce(i: %): Integer ==
i case Integer => i
userError "invalid conversion target type"

coerce(f: DoubleFloat): % ==
f pretend %
autoCoerce(f: %): DoubleFloat ==            -- used for hard coercion
f : DoubleFloat
coerce(f: %): DoubleFloat ==
f case DoubleFloat => f
userError "invalid conversion target type"
coerce(s: Symbol): % ==
s pretend %
autoCoerce(s: %): Symbol ==                 -- used for hard coercion
s : Symbol
coerce(s: %): Symbol ==
s case Symbol => s
userError "invalid conversion target type"
coerce(s: String): % ==
s pretend %
autoCoerce(s: %): String ==                 -- used for hard coercion
s : String
coerce(s: %): String ==
s case String => s
userError "invalid conversion target type"
buildSyntax(s: Symbol, l: List %): % ==
-- ??? ideally we should have overloaded operator `per' that convert
-- from list of syntax to syntax.  But the compiler is at the
-- moment defective for non-exported overloaded operations.
-- Furthermore, this direct call to `CONS' is currently necessary
-- in order to have the Syntax domain compiled as early as possible
-- in algebra boostrapping process.  It should be removed once
-- the bootstrap process is improved.
CONS(s,l)$Lisp @ %
buildSyntax(op: %, l: List %): % ==
CONS(op,l)$Lisp @ %
nil? x ==
null? rep x

getOperator x ==
atom? rep x => userError "atom as operand to getOperator"
op := car rep x
symbol? op => symbol op
integer? op => integer op
float? op => float op
string? op => string op
convert op
compound? x ==
pair? rep x
getOperands x ==
s := rep x
atom? s => []
[per t for t in destruct cdr s]

2 domain ConstructorCall

⟨domain CTORCALL ConstructorCall⟩≡

)abbrev domain CTORCALL ConstructorCall
++ Author: Gabriel Dos Reis
++ Date Created: January 19, 2008
++ Date Last Updated: January 19, 2008
++ Description: This domains represents a syntax object that
++ designates a category, domain, or a package.
++ See Also: Syntax, Domain
ConstructorCall(): Public == Private where
  Public ==> CoercibleTo OutputForm with
    constructorName: % -> Symbol
      ++ constructorName c returns the name of the constructor
    arguments: % -> List Syntax
      ++ arguments returns the list of syntax objects for the
      ++ arguments used to invoke the constructor.
  Private ==> add
    rep(x: %): List Syntax ==
      x pretend List(Syntax)
    constructorName x ==
      (first rep x)::Symbol
    arguments x ==
      rest rep x
    coerce x ==
      outputDomainConstructor(x)$Lisp

3 domain ElaboratedExpression

⟨domain ELABEXPR ElaboratedExpression⟩≡

)abbrev domain ELABEXPR ElaboratedExpression
++ Author: Gabriel Dos Reis
++ Date Created: January 19, 2008
++ Date Last Updated: January 20, 2008
++ Description: This domains an expresion as elaborated by the interpreter.
++ See Also:
ElaboratedExpression(): Public == Private where
Public ==> CoercibleTo OutputForm with
type: % -> ConstructorCall
++ type(e) returns the type of the expression as computed by
++ the interpreter.
constant?: % -> Boolean
++ constant?(e) returns true if `e' is a constant.
getConstant: % -> Union(SExpression,"failed")
++ getConstant(e) retrieves the constant value of `e'e.
variable?: % -> Boolean
++ variable?(e) returns true if `e' is a variable.
getVariable: % -> Union(Symbol,"failed")
++ getVariable(e) retrieves the name of the variable `e'.
callForm?: % -> Boolean
++ callForm?(e) is true when `e' is a call expression.
getOperator: % -> Union(Symbol, "failed")
++ getOperator(e) retrieves the operator being invoked in `e',
++ when `e' is an expression.
getOperands: % -> Union(List %, "failed")
++ getOperands(e) returns the of operands in `e'e, assuming it
++ is a call form.
Private ==> add
immediateDataTag := INTERN("--immediateData--")$Lisp
isAtomic(x: %): Boolean ==
ATOM(x)$Lisp @ Boolean
type x ==
getMode(x)$Lisp @ ConstructorCall
callForm? x ==
CONSP(x)$Lisp @ Boolean
getOperator x ==
op: SExpression := getUnnameIfCan(x)$Lisp
null? op => "failed"

op pretend Symbol
constant? x ==
isAtomic x and
  EQ(getUnnameIfCan(x)$Lisp, immediateDataTag)$Lisp : Boolean
getConstant x ==
constant? x => getValue(x)$Lisp @ SExpression
"failed"
variable? x ==
isAtomic x and not constant? x
getVariable x ==
variable? x => symbol (getUnname(x)$Lisp@SExpression)
"failed"

4 License

⟨license⟩≡

--Copyright (C) 2007, Gabriel Dos Reis.
--All rights reserved.
--
--Redistribution and use in source and binary forms, with or without
--modification, are permitted provided that the following conditions are
--met:
--
--    - Redistributions of source code must retain the above copyright
--      notice, this list of conditions and the following disclaimer.
--
--    - Redistributions in binary form must reproduce the above copyright
--      notice, this list of conditions and the following disclaimer in
--      the documentation and/or other materials provided with the
--      distribution.
--
--    - Neither the name of The Numerical Algorithms Group Ltd. nor the
--      names of its contributors may be used to endorse or promote products
--      derived from this software without specific prior written permission.
--
--THIS SOFTWARE IS PROVIDED BY THE COPYRIGHT HOLDERS AND CONTRIBUTORS "AS
--IS" AND ANY EXPRESS OR IMPLIED WARRANTIES, INCLUDING, BUT NOT LIMITED
--TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A
--PARTICULAR PURPOSE ARE DISCLAIMED. IN NO EVENT SHALL THE COPYRIGHT OWNER
--OR CONTRIBUTORS BE LIABLE FOR ANY DIRECT, INDIRECT, INCIDENTAL, SPECIAL,
--EXEMPLARY, OR CONSEQUENTIAL DAMAGES (INCLUDING, BUT NOT LIMITED TO,
--PROCUREMENT OF SUBSTITUTE GOODS OR SERVICES; LOSS OF USE, DATA, OR
--PROFITS; OR BUSINESS INTERRUPTION) HOWEVER CAUSED AND ON ANY THEORY OF
--LIABILITY, WHETHER IN CONTRACT, STRICT LIABILITY, OR TORT (INCLUDING
--NEGLIGENCE OR OTHERWISE) ARISING IN ANY WAY OUT OF THE USE OF THIS
--SOFTWARE, EVEN IF ADVISED OF THE POSSIBILITY OF SUCH DAMAGE.

⟨* ⟩≡ ⟨license⟩ ⟨domain SYNTAX Syntax⟩ ⟨domain CTORCALL ConstructorCall⟩ ⟨domain ELABEXPR ElaboratedExpression⟩